OpenStudy (anonymous):

Suppose you turn off the engine of your car when the temperature of the engine reaches 190 °F. If the outside temperature is a constant 65 °F, then the temperature T of the engine t minutes after you turn off the engine satisfies the equation below. ln T − 65 125 = −0.07t (a) Solve the equation for T. (b) Find the temperature of the engine 22 minutes after you turn it off. Give your answer to the nearest tenth of a degree.

OpenStudy (anonymous):

\[\ln((T-65)/125)=-0.07t\]

OpenStudy (anonymous):

first get rid of ln by e^(everything)

OpenStudy (anonymous):

(T-65)/125=ln(-0.07t

OpenStudy (anonymous):

or e^-0.07t

OpenStudy (anonymous):

\[(T-65)/125=e^{-0.07t}\]

OpenStudy (anonymous):

(T-65)/125 = e^(-0.07t) yea then times 125 both sides and then add 65 both sides to get T alone

OpenStudy (anonymous):

thank you?

OpenStudy (anonymous):

o: you re welcome you know how to do b, right?

OpenStudy (anonymous):

so: T = 125*e^(-0.07t) +65 (since t = 22 minutes) T = 125*e^(-0.07*22) + 65 T = 125*e^(-1.54) + 65 T = 125*0.2143811 + 65 T = 91.79763 ºF

OpenStudy (anonymous):

near to a tenth of a degree, T = 91.8 ºF